A divalent element x has a specific heat of 0.106 cal/g/°C. When a solution of the sulphate of x is electrolysed, 4.85g of x is deposited at the cathode and one litre of oxygen gas, measured at 16°C and 743mmHg pressure is evolved at the anode. Calculate the relative atomic mass of x (molar heat capacity =6.4 cal/mol/°C)
1
Expert's answer
2020-08-21T04:11:18-0400

reaction of electrolysis:

2XSO4 + 2H2O = 2X + 2H2SO4 + O2

n(O2) = pV/(RT) = 0.041 mol

n(X) = 2n(O2) = 0.082 mol

M(X) = m(X)/n(X) = 59 g mol-1